A person can see 74 miles to the horizon from an airplane window. How high is the airplane? Round to the nearest hundredth place. Use the formula V = 3.8√h, where h represents height and V is distance to the horizon.
Please tell me if my answer is right
v=3.8√h
74=3.8√h
h=(v/3.8 )^2=(74/3.)^2=379 ft

Answer by stanbon(75887)

A person can see 74 miles to the horizon from an airplane window.
How high is the airplane?
Round to the nearest hundredth place.
Use the formula V = 3.8√h, where h represents
height and V is distance to the horizon.
Please tell me if my answer is right
v=3.8√h
74=3.8√h
h=(v/3.8 )^2=(74/3.8)^2=379 ft

------------------
Correct
==========
Cheers,
Stan H.